# VERA: Variational Inference Framework for Jailbreaking Large Language Models

### v1

@@ Abstract
The rise of API-only access to state-of-the-art LLMs highlights the need for effective black-box jailbreak methods to identify model vulnerabilities in real-world settings. Without a principled objective for gradient-based optimization, most existing approaches rely on genetic algorithms, which are limited by their initialization and dependence on manually curated prompt pools. Furthermore, these methods require individual optimization for each prompt, failing to provide a comprehensive characterization of model vulnerabilities. To address this gap, we introduce VERA: Variational infErence fRamework for jAilbreaking. VERA casts black-box jailbreak prompting as a variational inference problem, training a small attacker LLM to approximate the target LLM’s posterior over adversarial prompts. Once trained, the attacker can generate diverse, fluent jailbreak prompts for a target query without re-optimization. Experimental results show that VERA achieves strong performance across a range of target LLMs, highlighting the value of probabilistic inference for adversarial prompt generation. The code will be made available upon publication.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed across various applications due to their impressive capabilities. However, these models remain vulnerable to jailbreaking attacks, where adversaries craft prompts to bypass safety guardrails and elicit harmful responses [43]. Effective red-teaming methods are crucial for conducting adversarial testing and identifying failure modes.
Existing jailbreaking approaches fall into several categories. Some works focus on manually crafted prompts that expose the vulnerability of LLM safeguarding techniques [33, 27]. While these methods reveal previously unknown vulnerabilities in LLMs, they lack scalability and coverage, as constructing such prompts requires substantial human effort and domain expertise.
To address these limitations, a growing body of work focuses on automated adversarial prompt design. Such works can be categorized based on how much access the adversary has to the target LLM. White-box attacks assume that the attacker has access to the entire target LLM, which enables the use of gradient-based optimization and sampling [7, 43, 9, 42]. Black-box attacks assume the adversary only has API access to the target LLM, reflecting the threat model faced during deployment. These methods rely on gradient-free optimization techniques, such as genetic algorithms [18, 39, 15] or search algorithms [3, 20, 16].
Despite their effectiveness, existing automated black-box attack methods have notable drawbacks. They require individual optimization or search loops for each prompt, resulting in high computational costs [40, 18, 3]. This makes them impractical for comprehensive red teaming, which requires characterizing the breadth of model vulnerability for a specific behavior, rather than identifying isolated instances of failure. Moreover, they often depend on initialization from pools of manually crafted prompts that are known to work [18, 39, 26, 41]. This ties these methods to specific vulnerabilities that are already known and therefore more likely to be addressed by model developers. As a result, current jailbreak methods are fragile, as they rely on prompt patterns that are likely to be detected, patched, or rendered obsolete as alignment techniques improve.
@@
Figure 1: Comparison between traditional jailbreak pipelines and VERA. Traditional methods (e.g., [3, 39, 18]) require per-prompt mutation, scoring, and iterative querying to generate a successful jailbreak, making them slow and brittle. In contrast, VERA samples diverse, high-quality adversarial prompts directly from a learned distribution at test time, enabling fast, parallelizable prompt generation without any additional optimization or search.
@@ 2 Related work
Jailbreaking large language models (LLMs) refers to crafting inputs that elicit restricted or unsafe outputs, despite safety filters. Early work explored manual prompt engineering [3, 28], which lacked scalability and robustness against patching. More recent efforts leverage optimization-based methods to automate jailbreak discovery.
White Box Attacks. White-box approaches exploit gradient access to perturb prompts for adversarial success. Greedy Coordinate Gradient (GCG) and its variants [7, 14, 2] iteratively update tokens to maximize harmful output likelihood, but require extensive model access and produce brittle, incoherent prompts. Subsequent work addresses fluency constraints through left-to-right decoding [42] or controlled text generation [9], yielding more readable outputs. However, these methods still suffer from fixed token commitments and brittle multi-objective balancing, limiting their adaptability and efficiency.
Black Box Attacks. Black-box methods remove the reliance on model internals, aligning more realistically with API-only settings. Strategies like genetic algorithms, such as GPTFuzzer [39], AutoDAN [18], evolve prompts through random mutation and selection, which often leads to inefficient exploration due to their stochastic and undirected nature. LLM-guided techniques like PAIR [3] and TAP [20] restructure queries based on model feedback. While these prompt-based methods improve interpretability, they lack clear optimization guidance for the jailbreak objective and suffer from limited diversity in generated prompts. More recent approaches use reinforcement learning (RL) agents [4, 5] to perform a more guided search than random genetic algorithm mutations. However, they are not end-to-end and have a two-stage optimization problem that separates strategy selection from prompt generation.
LLM Defenses. Recent defense strategies aim to mitigate adversarial attacks. Baseline defenses include perplexity-based detection, input preprocessing, and adversarial training [12]. More advanced techniques, such as Circuit Breakers, directly control the representations responsible for harmful outputs [44]. Additionally, Llama Guard employs LLM-based classifiers to filter unsafe inputs and outputs [11]. While these defenses can reduce attack success rates, they are not foolproof and remain vulnerable to adaptive adversaries.
@@ 3.1 Variational framework
Variational objective. To solve this problem, we use a pretrained LLM, referred to as the attacker LLM, as the variational distribution $q_{\theta}(x)$ to approximate the posterior distribution over adversarial prompts. We parameterize $q_{\theta}(x)$ using LoRA adaptors as it makes fine-tuning relatively cheap [10]. In this case, $\theta$ denotes the LoRA parameters. We define the variational objective as follows:
@@
$P_{LM}(y^{*}|x)\approx J(x,\hat{y}),$ (5)
where $\hat{y}$ is the response produced from the prompt $x$. Intuitively, we approximate the probability of a given prompt $x$ generating a harmful response $y\in\mathcal{Y}_{\text{harm}}$ by using the normalized Judge score, which outputs a scalar in $[0,1]$ that measures the harmfulness of the response $\hat{y}$. In practice, the Judge can be instantiated either as a lightweight binary classifier [39, 19] or via prompting an LLM to produce a safety judgment score. When using a binary classifier, we interpret the softmax confidence of the positive (harmful) class as a probabilistic proxy. This flexibility allows us to extract a smooth guidance signal, facilitating gradient-based optimization.
REINFORCE gradient estimator. Directly optimizing the objective is difficult, as it requires taking a gradient of an expectation that depends on the target parameters $\theta$. To address this issue, we use the REINFORCE gradient estimator [37]. We define the function $f$ as follows:
@@ 3.2 VERA algorithm
[…]
3:  cur-best-val $\leftarrow$-$\infty$ - ∞
@@
7:        $x\sim q_{\theta}(\cdot)$triangleright These operations are straight forward to implement as batch computations on GPUs
@@
[…]
@@ 3.3 Advantages of variational jailbreaking
To demonstrate these advantages, we sample a subset of 50 behaviors from the HarmBench dataset and experiment against Vicuna 7B as the target LLM. We compare against GPTFuzzer and AutoDAN [^1] We use the Harmbench implementations for both methods., as representative template-based automated black-box algorithms. These utilize human-written jailbreak templates as initial seeds and iteratively generate new prompts. Since many of the templates GPTFuzzer and AutoDAN use as initialization can jailbreak the target without any edits, we also show the performance when both methods are restricted to templates that are incapable of jailbreaking the target on their own [39, 18]. We label these versions with an asterisk. We assess performance under two evaluation protocols:
@@
Independence from manual templates. Many automated black-box attack methods bootstrap the discovery of effective prompts from manually crafted prompts that are known to be effective [39]. While this has been shown to improve ASR, this also renders such methods dependent on known vulnerabilities, which are the easiest to patch from a model developer’s perspective. Thus, it is desirable to have jailbreaking methods that are fully autonomous and independent of initial templates.
To demonstrate that VERA is independent of the initial system prompt, we compare the similarity of the generated attacks per behavior with the system prompt (Figure  5) by using the BLEU score. When computing this metric for GPTFuzzer and AutoDAN, we use their set of initial templates to be the reference texts.
@@ 4.1 Experimental setup
Dataset. We evaluate our approach on the HarmBench dataset [19], a comprehensive benchmark for evaluating jailbreak attacks and robust refusal in LLMs. HarmBench consists of 400 harmful behaviors curated with reference to content policies, spanning 7 diverse categories such as illegal activities, hate speech, and misinformation. This dataset has become a standard evaluation framework for automated red teaming and serves as a robust benchmark for assessing attack effectiveness.
Target models. We evaluate our method against 8 large language models, including 6 open-source models and 2 commercial models, chosen from the HarmBench leaderboard [19], representing state-of-the-art results in jailbreaking research. Our open-source targets include LLaMA2-7b-chat and LLaMA2-13b-chat [31], Vicuna-7b [6], Baichuan-2-7b [38], Orca2-7b [21], and Zephyr-7b-robust (adversarially trained Zephyr-7b using Robust Refusal Dynamic Defense against a GCG adversary)  [32, 19]. For proprietary models, we evaluate on Gemini-Pro [30] and GPT-3.5-Turbo1106 [22].
Attacker models. We use Vicuna-7b chat as our default attacker model, consistent with widespread adoption in current literature [3, 20, 35] due to its strong compliance capabilities. To demonstrate the generalizability of VERA, we include an ablation study in Section A.2 examining performance across different attacker model architectures.
Evaluation metrics. Following established protocols in jailbreaking research [17, 43, 19], we measure Attack Success Rate (ASR), defined as the percentage of prompts that successfully induce the target model to generate harmful content complying with the malicious instruction. Success determination follows the HarmBench protocol, utilizing a fine-tuned LLaMA2-13B classifier.
Baselines. We compare VERA against a comprehensive set of jailbreaking approaches from the HarmBench leaderboard. These include gradient-based white-box methods such as GCG [43], GCG-M [43] and GCG-T [43], PEZ [36], GBDA [8], UAT [34], and AP [29]; black-box techniques including SFS [24]; ZS [24], PAIR [3], TAP and TAP-T [20]; evolutionary algorithms such as AutoDAN [18] and PAP-top5 [39]; and Human Direct [28], representing manually crafted jailbreak prompts. We exclude AutoDAN-Turbo [17] from our comparison as we were unable to reproduce their results in our experimental environment. For further details on experimental design and the hardware used to run these experiments, refer to Appendix B.
@@ 4.2 Main results
Table 1: Our method VERA is the state-of-the-art attack in Harmbench [19]. The upper section of the table lists white-box baselines, while the lower section lists black-box baselines. Bold values indicate the best performance across all methods, and underlined values highlight the best among black-box approaches.
@@
We include the main results for the HarmBench benchmark in Table 1. Our method achieves state-of-the-art performance across open-source models, outperforming all prior black-box and white-box approaches. Notably, VERA attains an ASR of 70.0% on Vicuna-7B, 64.8% on Baichuan2-7B, 72.0% on Orca2-7B, 63.5% on R2D2, and 48.5% on Gemini-Pro, outperforming GCG, AutoDAN, TAP-T, and other methods. This demonstrates our method’s ability to craft highly effective adversarial prompts that generalize across model architectures. Example generations are provided in Appendix C.
Although gradient-based GCG and its variants utilize full access to model internals, VERA outperforms them on five of the seven open-source targets, trailing only on the LLaMA2 family. This gap can be attributed to LLaMA2’s strong RLHF-based safety alignment. Unlike white-box methods that can exploit LLaMA’s internal structure, VERA operates purely in the black-box setting, relying only on output signals, making the optimization problem significantly harder for models with robust and deterministic safety filters. Nevertheless, VERA still outperforms all black-box baselines on the LLaMA2 models, highlighting its effectiveness even when the target LLM gradients and internals are inaccessible.
@@ 4.4 Robustness to defenses
To evaluate the practical viability of VERA, we assess its performance against two widely adopted defense mechanisms: the Perplexity Filter (PF)[13, 1], which blocks prompts deemed incoherent or low-likelihood under a language model, and Circuit Breaker [44] (CB), a defense technique that monitors internal activations to interrupt harmful generations.  [^2] We use the official GitHub implementations for both defenses. For CB, we follow the available configuration for LLaMA3-8B-Instruct, as the official repository provides support only for LLaMA3-8B-Instruct and Mistral.
@@
These results demonstrate that VERA produces adversarial prompts that are robust to both static and dynamic defenses, raising a new concern for the LLM safety community.

### v2

@@ Abstract
The rise of API-only access to state-of-the-art LLMs highlights the need for effective black-box jailbreak methods to identify model vulnerabilities in real-world settings. Without a principled objective for gradient-based optimization, most existing approaches rely on genetic algorithms, which are limited by their initialization and dependence on manually curated prompt pools. Furthermore, these methods require individual optimization for each prompt, failing to provide a comprehensive characterization of model vulnerabilities. To address this gap, we introduce VERA: Variational infErence fRamework for jAilbreaking. VERA casts black-box jailbreak prompting as a variational inference problem, training a small attacker LLM to approximate the target LLM’s posterior over adversarial prompts. Once trained, the attacker can generate diverse, fluent jailbreak prompts for a target query without re-optimization. Experimental results show that VERA achieves strong performance across a range of target LLMs, highlighting the value of probabilistic inference for adversarial prompt generation.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed across various applications due to their impressive capabilities. However, these models remain vulnerable to jailbreaking attacks, where adversaries craft prompts to bypass safety guardrails and elicit harmful responses [zou2023universal]. Effective red-teaming methods are crucial for conducting adversarial testing and identifying failure modes.
Existing jailbreaking approaches fall into several categories. Some works focus on manually crafted prompts that expose the vulnerability of LLM safeguarding techniques [wallace-etal-2019-universal, shen2023do]. While these methods reveal previously unknown vulnerabilities in LLMs, they lack scalability and coverage, as constructing such prompts requires substantial human effort and domain expertise.
To address these limitations, a growing body of work focuses on automated adversarial prompt design. Such works can be categorized based on how much access the adversary has to the target LLM. White-box attacks assume that the attacker has access to the entire target LLM, which enables the use of gradient-based optimization and sampling [guo-etal-2021-gradient, zou2023universal, guo2024coldattack, zhu2023autodan]. Black-box attacks assume the adversary only has API access to the target LLM, reflecting the threat model faced during deployment. These methods rely on gradient-free optimization techniques, such as genetic algorithms [liu2024autodan, yu2023gptfuzzer, lapid2023open] or search algorithms [chao2023jailbreaking, mehrotra2024tree, li2024semantic].
Despite their effectiveness, existing automated black-box attack methods have notable drawbacks. They require individual optimization or search loops for each prompt, resulting in high computational costs [yuan2024gpt4, liu2024autodan, chao2023jailbreaking]. This makes them impractical for comprehensive red teaming, which requires characterizing the breadth of model vulnerability for a specific behavior, rather than identifying isolated instances of failure. Moreover, they often depend on initialization from pools of manually crafted prompts that are known to work [liu2024autodan, yu2023gptfuzzer, shah2023scalable, zeng2024johnny]. This ties these methods to specific vulnerabilities that are already known and therefore more likely to be addressed by model developers. As a result, current jailbreak methods are fragile, as they rely on prompt patterns that are likely to be detected, patched, or rendered obsolete as alignment techniques improve.
@@
Figure 1: Comparison between traditional jailbreak pipelines and VERA. Traditional methods (e.g., [chao2023jailbreaking, yu2023gptfuzzer, liu2024autodan]) require per-prompt mutation, scoring, and iterative querying to generate a successful jailbreak, making them slow and brittle. In contrast, VERA samples diverse, high-quality adversarial prompts directly from a learned distribution at test time, enabling fast, parallelizable prompt generation without any additional optimization or search.
@@ 2 Related work
Jailbreaking large language models (LLMs) refers to crafting inputs that elicit restricted or unsafe outputs, despite safety filters. Early work explored manual prompt engineering [chao2023jailbreaking, shen2024anything], which lacked scalability and robustness against patching. More recent efforts leverage optimization-based methods to automate jailbreak discovery.
White Box Attacks. White-box approaches exploit gradient access to perturb prompts for adversarial success. Greedy Coordinate Gradient (GCG) and its variants [guo-etal-2021-gradient, jia2024improvedtechniquesoptimizationbasedjailbreaking, anonymous2024ebgcg] iteratively update tokens to maximize harmful output likelihood, but require extensive model access and produce brittle, incoherent prompts. Subsequent work addresses fluency constraints through left-to-right decoding [zhu2023autodan] or controlled text generation [guo2024coldattack], yielding more readable outputs. However, these methods still suffer from fixed token commitments and brittle multi-objective balancing, limiting their adaptability and efficiency.
Black Box Attacks. Black-box methods remove the reliance on model internals, aligning more realistically with API-only settings. Strategies like genetic algorithms, such as GPTFuzzer [yu2023gptfuzzer], AutoDAN [liu2024autodan], evolve prompts through random mutation and selection, which often leads to inefficient exploration due to their stochastic and undirected nature. LLM-guided techniques like PAIR [chao2023jailbreaking] and TAP [mehrotra2024tree] restructure queries based on model feedback. While these prompt-based methods improve interpretability, they lack clear optimization guidance for the jailbreak objective and suffer from limited diversity in generated prompts. More recent approaches use reinforcement learning (RL) agents [chen2024llm, chen2024rl, rlta] to perform a more guided search than random genetic algorithm mutations. However, they are not end-to-end and have a two-stage optimization problem that separates strategy selection from prompt generation. In rlta authors fine-tune a language model as a general-purpose attacker, but incur substantial training overhead (96 hours) and often default to generic exploits, missing behavior-specific failure modes.
LLM Defenses. Recent defense strategies aim to mitigate adversarial attacks. Baseline defenses include perplexity-based detection, input preprocessing, and adversarial training [jain2023baseline]. More advanced techniques, such as Circuit Breakers, directly control the representations responsible for harmful outputs [zou2024improving]. Additionally, Llama Guard employs LLM-based classifiers to filter unsafe inputs and outputs [inan2023llama]. While these defenses can reduce attack success rates, they are not foolproof and remain vulnerable to adaptive adversaries.
@@ 3.1 Variational framework
Variational objective. To solve this problem, we use a pretrained LLM, referred to as the attacker LLM, as the variational distribution $q_{\theta}(x)$ to approximate the posterior distribution over adversarial prompts. We parameterize $q_{\theta}(x)$ using LoRA adaptors as it makes fine-tuning relatively cheap [hu2022lora]. In this case, $\theta$ denotes the LoRA parameters. We define the variational objective as follows:
@@
[…]
where $\hat{y}$ is the response produced from the prompt $x$. Intuitively, we approximate the probability of a given prompt $x$ generating a harmful response $y\in\mathcal{Y}_{\text{harm}}$ by using the normalized Judge score, which outputs a scalar in $[0,1]$ that measures the harmfulness of the response $\hat{y}$. In practice, the Judge can be instantiated either as a lightweight binary classifier [yu2023gptfuzzer, mazeika2024harmbench] or via prompting an LLM to produce a safety judgment score. When using a binary classifier, we interpret the softmax confidence of the positive (harmful) class as a probabilistic proxy. This flexibility allows us to extract a smooth guidance signal, facilitating gradient-based optimization.
REINFORCE gradient estimator. Directly optimizing the objective is difficult, as it requires taking a gradient of an expectation that depends on the target parameters $\theta$. To address this issue, we use the REINFORCE gradient estimator [williams1992simple]. We define the function $f$ as follows:
@@ 3.2 VERA algorithm
2: cur-best $\leftarrow\emptyset$
3: cur-best-val $\leftarrow$-$\infty$
@@
7:   $x\sim q_{\theta}(\cdot)$ $\triangleright$ These operations are straight forward to implement as batch computations on GPUs
@@
16:  $\nabla_{\theta}ELBO(\text{cur-prompt, cur-resp, cur-scores})\leftarrow$ compute REINFORCE estimator using (8).
17:  $\theta\leftarrow\theta+\gamma\nabla_{\theta}ELBO(\text{cur-prompt, cur-resp, cur-scores})$
@@ 3.3 Advantages of variational jailbreaking
To demonstrate these advantages, we sample a subset of 50 behaviors from the HarmBench dataset and experiment against Vicuna 7B as the target LLM. We compare against GPTFuzzer and AutoDAN [^1] We use the Harmbench implementations for both methods., as representative template-based automated black-box algorithms. These utilize human-written jailbreak templates as initial seeds and iteratively generate new prompts. Since many of the templates GPTFuzzer and AutoDAN use as initialization can jailbreak the target without any edits, we also show the performance when both methods are restricted to templates that are incapable of jailbreaking the target on their own [yu2023gptfuzzer, liu2024autodan]. We label these versions with an asterisk. We assess performance under two evaluation protocols:
@@
Independence from manual templates. Many automated black-box attack methods bootstrap the discovery of effective prompts from manually crafted prompts that are known to be effective [yu2023gptfuzzer]. While this has been shown to improve ASR, this also renders such methods dependent on known vulnerabilities, which are the easiest to patch from a model developer’s perspective. Thus, it is desirable to have jailbreaking methods that are fully autonomous and independent of initial templates.
To demonstrate that VERA is independent of the initial system prompt, we compare the similarity of the generated attacks per behavior with the system prompt (Figure 5) by using the BLEU score. When computing this metric for GPTFuzzer and AutoDAN, we use their set of initial templates to be the reference texts.
@@ 4.1 Experimental setup
Dataset. We evaluate our approach on the HarmBench dataset [mazeika2024harmbench], a comprehensive benchmark for evaluating jailbreak attacks and robust refusal in LLMs. HarmBench consists of 400 harmful behaviors curated with reference to content policies, spanning 7 diverse categories such as illegal activities, hate speech, and misinformation. This dataset has become a standard evaluation framework for automated red teaming and serves as a robust benchmark for assessing attack effectiveness.
Target models. We evaluate our method against 8 large language models, including 6 open-source models and 2 commercial models, chosen from the HarmBench leaderboard [mazeika2024harmbench], representing state-of-the-art results in jailbreaking research. Our open-source targets include LLaMA2-7b-chat and LLaMA2-13b-chat [touvron2023llama], Vicuna-7b [vicuna2023], Baichuan-2-7b [yang2025baichuan2openlargescale], Orca2-7b [mitra2023orca], and Zephyr-7b-robust (adversarially trained Zephyr-7b using Robust Refusal Dynamic Defense against a GCG adversary) [tunstall2023zephyr, mazeika2024harmbench]. For proprietary models, we evaluate on Gemini-Pro [team2023gemini] and GPT-3.5-Turbo1106 [openai2024gpt4technicalreport].
Attacker models. We use Vicuna-7b chat as our default attacker model, consistent with widespread adoption in current literature [chao2023jailbreaking, mehrotra2024tree, wei2023jailbroken] due to its strong compliance capabilities. To demonstrate the generalizability of VERA, we include an ablation study in Section LABEL:sec:abl examining performance across different attacker model architectures.
Evaluation metrics. Following established protocols in jailbreaking research [liu2024autodanturbo, zou2023universal, mazeika2024harmbench], we measure Attack Success Rate (ASR), defined as the percentage of prompts that successfully induce the target model to generate harmful content complying with the malicious instruction. Success determination follows the HarmBench protocol, utilizing a fine-tuned LLaMA2-13B classifier.
Baselines. We compare VERA against a comprehensive set of jailbreaking approaches from the HarmBench leaderboard. These include gradient-based white-box methods such as GCG [zou2023universal], GCG-M [zou2023universal] and GCG-T [zou2023universal], PEZ [wen2023hard], GBDA [guo2021gradient], UAT [wallace2019universal], and AP [shin2020autoprompt]; black-box techniques including SFS [perez2022red]; ZS [perez2022red], PAIR [chao2023jailbreaking], TAP and TAP-T [mehrotra2024tree]; evolutionary algorithms such as AutoDAN [liu2024autodan] and PAP-top5 [yu2023gptfuzzer]; and Human Direct [shen2024anything], representing manually crafted jailbreak prompts. We exclude AutoDAN-Turbo [liu2024autodanturbo] from our comparison as we were unable to reproduce their results in our experimental environment. For further details on experimental design and the hardware used to run these experiments, refer to Appendix D.
@@ 4.2 Main results
Table 1: Our method VERA is the state-of-the-art attack in Harmbench [mazeika2024harmbench]. The upper section of the table lists white-box baselines, while the lower section lists black-box baselines. Bold values indicate the best performance across all methods, and underlined values highlight the best among black-box approaches.
@@
We include the main results for the HarmBench benchmark in Table 1. Our method achieves state-of-the-art performance across open-source models, outperforming all prior black-box and white-box approaches. Notably, VERA attains an ASR of 70.0% on Vicuna-7B, 64.8% on Baichuan2-7B, 72.0% on Orca2-7B, 63.5% on R2D2, and 48.5% on Gemini-Pro, outperforming GCG, AutoDAN, TAP-T, and other methods. This demonstrates our method’s ability to craft highly effective adversarial prompts that generalize across model architectures. Example generations are provided in Appendix E.
Although gradient-based GCG and its variants utilize full access to model internals, VERA outperforms them on five of the seven open-source targets, trailing only on the LLaMA2 family. This gap can be attributed to LLaMA2’s strong RLHF-based safety alignment. Unlike white-box methods that can exploit LLaMA’s internal structure, VERA operates purely in the black-box setting, relying only on output signals, making the optimization problem significantly harder for models with robust and deterministic safety filters. Nevertheless, VERA still outperforms all black-box baselines on the LLaMA2 models, highlighting its effectiveness even when the target LLM gradients and internals are inaccessible. In addition to HarmBench [mazeika2024harmbench], we evaluate VERA on the AdvBench benchmark (see Appendix C), showing similar trends in performance.
@@ 4.4 Robustness to defenses
To evaluate the practical viability of VERA, we assess its performance against two widely adopted defense mechanisms: the Perplexity Filter (PF)[jain2023baselinedefensesadversarialattacks, alon2023detectinglanguagemodelattacks], which blocks prompts deemed incoherent or low-likelihood under a language model, and Circuit Breaker [zou2024improving] (CB), a defense technique that monitors internal activations to interrupt harmful generations.  [^2] We use the official GitHub implementations for both defenses. For CB, we follow the available configuration for LLaMA3-8B-Instruct, as the official repository provides support only for LLaMA3-8B-Instruct and Mistral.
@@
These results demonstrate that VERA produces adversarial prompts that are robust to both static and dynamic defenses, raising a new concern for the LLM safety community. We further evaluate VERA under recently proposed defense mechanisms LLaMA Guard, SmoothLLM, and RA-LLM in Appendix C.
4.5 Ablation Studies
To analyse VERA’s performance, we conduct a series of ablation studies, detailed in Appendix B, examining the role of optimization (via comparison to Best-of-N baseline), the effect of different attacker LLM backbones, the impact of KL regularization coefficients, and the influence of judge model choice on training outcomes.